A basic cellular package costs $30/month for 60 minutes of calling with an additional charge of $0.50/minute beyond that time. The cost function C(x) for using x minutes would be
If you used 60 minutes or less, i.e. if if x≤60, then C(x)=30 (the base charge).
If you used more than 60 minutes, i.e. (x−60) minutes more than the plan came with, you would pay an additional $0.50 for each of those (x−60) minutes. Your total bill would be C(x)=30+0.50(x−60).

If you want to keep your bill at $50 or lower for the month, what is the maximum number of calling minutes you can use?

Answer by Boreal(15235)

C(x)=30+0.50(x-60)<=50
0.50(x-60)<=20
0.5x-30<=20
0.5x < =50
x <= 100 minutes
100 minutes is the answer.